OpenStudy (anonymous):

simplify the expression √(z11)/(81z). The variable represents a nonnegative number.

OpenStudy (anonymous):

I think it would be something like this : \[\sqrt{\frac{ 11z }{ 81z }} = \sqrt{\frac{ 11 }{ 81 }} = + \frac{ \sqrt{11} }{ 9 }\]

OpenStudy (anonymous):